5. (6 points) Analyze the graph of the rational function f(x) = x + x -12 x² - 4 Label the intercepts, asymptotes, if any.
Answer #1

g Ques) f(c) = 1²+2-12 ay Sofy (1) x-intercept: means y=0 o= x²+2-12 x² - 4 7+424-34-)2=0 H(x+4) – 3(x+4)=0 31+4) (-3)=0 n=-4